OpenStudy (ckallerid):

Given f(x) = the quantity of 2x minus 2 divided by 4, solve for f−1(3).

OpenStudy (amtran_bus):

Dude spongebob looks awful

OpenStudy (ckallerid):

correction: sexy

OpenStudy (ckallerid):

\[f(x)=\frac{ 2x-2 }{ 4 }\]

OpenStudy (ckallerid):

except needs to be inverse

OpenStudy (lynfran):

find the inverse of f(x) first f(x)=(2x-2)/4 y=(2x-2)/4 solve for x we get 4y=2x-2 4y+2=2x (4y+2)/2=x replace x and y we get its inverse (4x+2)/2= y f-1(x)=(4x+2)/2 f-1(3)=?? sub 3 where x is to solve...

OpenStudy (ckallerid):

@LynFran So it's f(3) = 7?

OpenStudy (lynfran):

yep

OpenStudy (ckallerid):

okie thnx! :D